in hydrolysis we do rxn with h2o or h3o???????????

Grade:12

1 Answers

AskiitianExpert Pramod-IIT-R
14 years ago

Hydrolysis

IT  is a chemical reaction during which molecules of water (H2O) are split into hydrogen (H) and hydroxide anions (OH) in the process of a chemical mechanism. It is the type of reaction that is used to break down certain polymers, especially those made by step-growth polymerization. Such polymer degradation is usually catalysed by either acid, e.g., concentrated sulfuric acid2SO4), or alkali, e.g., sodium hydroxide (NaOH) attack, often increasing with their strength or pH. (H

Hydrolysis is distinct from hydration. In hydration, the hydrated molecule does not "lyse" (break into two new compounds). It should not be confused with hydrogenolysis, a reaction of hydrogen.

Contents

Types

Hydrolysis is a chemical process in which a molecule is cleaved into two parts by the addition of a molecule of water. One fragment of the parent molecule gains a hydrogen ion (H+) from the additional water molecule. The other group collects the remaining hydroxyl group (OH).

The most common hydrolysis occurs when a salt of a weak acid or weak base (or both) is dissolved in water. Water autoionizes into negative hydroxyl ions and hydrogen ions. The salt breaks down into positive and negative ions. For example, sodium acetateacetate ions. Sodium ions react very little with hydroxyl ions whereas acetate ions combine with hydrogen ions to produce neutral acetic acid, and the net result is a relative excess of hydroxyl ions, causing a basic solution. dissociates in water into sodium and acetate ions.

However, under normal conditions, only a few reactions between water and organic compounds occur. In general, strong acids or bases must be added in order to achieve hydrolysis where water has no effect. The acid or base is considered a catalyst. They are meant to speed up the reaction, but are recovered at the end of it.

Acid–base-catalyzed hydrolyses are very common; one example is the hydrolysis of amides or esters. Their hydrolysis occurs when the nucleophile (a nucleus-seeking agent, e.g., water or hydroxyl ion) attacks the carbon of the carbonyl group of the ester or amide. In an aqueous base, hydroxyl ions are better nucleophiles than dipoles such as water. In acid, the carbonyl group becomes protonated, and this leads to a much easier nucleophilic attack. The products for both hydrolyses are compounds with carboxylic acid groups.

Perhaps the oldest example of ester hydrolysis is the process called saponification. It is the hydrolysis of a triglyceride (fat) with an aqueous base such as sodium hydroxide (NaOH). During the process, glycerol is formed, and the fatty acids react with the base, converting them to salts. These salts are called soaps, commonly used in households.

Moreover, hydrolysis is an important process in plants and animals, the most significant example being energy metabolism and storage. All living cells require a continual supply of energy for two main purposes: for the biosynthesis of small and macromolecules, and for the active transport of ions and molecules across cell membranes. The energy derived from the oxidationadenosine triphosphate (ATP). of nutrients is not used directly but, by means of a complex and long sequence of reactions, it is channeled into a special energy-storage molecule,

The ATP molecule contains pyrophosphate linkages (bonds formed when two phosphate units are combined together) that release energy when needed. ATP can undergo hydrolysis in two ways: the removal of terminal phosphate to form adenosine diphosphateadenosine monophosphate (AMP) and pyrophosphate. The latter is usually cleaved further to yield two phosphates. This results in biosynthesis reactions, which do not occur alone, that can be driven in the direction of synthesis when the phosphate bonds have undergone hydrolysis. (ADP) and inorganic phosphate, or the removal of a terminal diphosphate to yield

In addition, in living systems, most biochemical reactions, including ATP hydrolysis, take place during the catalysis of enzymes. The catalytic action of enzymes allows the hydrolysis of proteins, fats, oils, and carbohydrates. As an example, one may consider proteases, enzymes that aid digestion by causing hydrolysis of peptide bonds in proteins. They catalyze the hydrolysis of interior peptide bonds in peptide chains, as opposed to exopeptidases, another class of enzymes, that catalyze the hydrolysis of terminal peptide bonds, liberating one free amino acid at a time.

However, proteases do not catalyze the hydrolysis of all kinds of proteins. Their action is stereo-selective: Only proteins with a certain tertiary structure will be targeted. The reason is that some kind of orienting force is needed to place the amide group in the proper position for catalysis. The necessary contacts between an enzyme and its substrates (proteins) are created because the enzyme folds in such a way as to form a crevice into which the substrate fits; the crevice also contains the catalytic groups. Therefore, proteins that do not fit into the crevice will not undergo hydrolysis. This specificity preserves the integrity of other proteins such as hormones, and therefore the biological system continues to function normally.

[edit] Hydrolysis of amide links

In the hydrolysis of an amide into a carboxylic acid and an amine or ammonia, the carboxylic acid has a hydroxyl group derived from a water molecule and the amine (or ammonia) gains the hydrogen ion.

Amide hydrolysis.png

A specific case of the hydrolysis of an amide link is the hydrolysis of peptides to smaller fragments or amino acids.

Many polyamide polymers such as nylon 6,6 are attacked and hydrolysed in the presence of strong acids. Such attack leads to depolymerization and nylon products fail by fracturing when exposed to even small amounts of acid. Other polymers made by step-growth polymerization are susceptible to similar polymer degradation reactions. The problem is known as stress corrosion cracking.

[edit] Hydrolysis of polysaccharides

Sucrose. The glycoside bond is represented by the central oxygen atom, which holds the two monosaccharide units together.

Monosaccharides can be linked together by glycosidic bonds, which can be cleaved by hydrolysis. Two, three, several or many monosaccharides thus linked form disaccharides, trisaccharides, oligosaccharides or polysaccharides, respectively. Enzymes that hydrolyse glycosidic bonds are called "glycoside hydrolases" or "glycosidases".

The best-known disaccharide is sucrose (table sugar). Hydrolysis of sucrose yields glucose and fructose. Invertase is a sucraseinvert sugar. Lactase is essential for digestive hydrolysis of lactose in milk. Deficiency of lactase in humans causes lactose intolerance. used industrially for the hydrolysis of sucrose to so-called

The hydrolysis of polysaccharides to soluble sugars is called "saccharification". Malt made from barley is used as a source of β-amylase to break down starch into the disaccharide maltose, which can be used by yeast to produce beer. Other amylaseCellulose is converted to glucose or the disaccharide cellobiose by cellulases. Animals such as cows (ruminants) are able to digest cellulose because of symbiotic bacteria that produce cellulases. enzymes may convert starch to glucose or to oligosaccharides.

[edit] Irreversibility of hydrolysis under physiological conditions

Under physiological conditions (e.g., in dilute aqueous solution), a hydrolytic cleavage reaction, in which the concentration of a metabolic precursor is low (on the order of 10−3 to 10−6 molar), is essentially thermodynamically irreversible. To give an example:

A + H2O → X + Y
K_d = \frac{\left[X\right] \left[Y\right]} {\left[H_2O\right] \left[A\right]}

Assuming that x is the final concentration of products, and that C is the initial concentration of A, and W = [H2O] = 55.5 molar, then x can be calculated with the equation:

\frac{x \times x}{W\left(C - x\right)} = K_d

let Kd×W = k:

then  x = \frac {-k + \sqrt {k^2 + 4kC} } {2}.

For a value of C = 0.001 molar, and k = 1 molar, x/C > 0.999. Less than 0.1% of the original reactant would be present once the reaction is complete.

This theme of physiological irreversibility of hydrolysis is used consistently in metabolic pathways, since many biological processes are driven by the cleavage of anhydrous pyrophosphate bonds.

[edit] Hydrolysis of metal aqua ions

Metal ions are Lewis acids, and in aqueous solution they form aqua ions, of the general formula M(H2O)nm+. [3] [4] The aqua ions undergo hydrolysis, to a greater or lesser extent. The first hydrolysis step is given generically as

M(H2O)nm+ + H2O is in equilibrium with M(H2O)n-1(OH)(m-1)+ + H3O+

Thus the aqua ion is behaving as an acid in terms of Brønsted-Lowry acid-base theory. This is easily explained by considering the inductive effect of the positively charged metal ion, which weakens the O-H bond of an attached water molecule, making the liberation of a proton relatively easy.

The dissociation constant, pKa, for this reaction is more or less linearly related to the charge-to-size ratio of the metal ion.[5] Ions with low charges, such as Na+ are very weak acids with almost imperceptible hydrolysis. Large divalent ions such as Ca2+, Zn2+, Sn2+ and Pb2+ have a pKa of 6 or more and would not normally be classed as acids, but small divalent ions such as Be2+ undergo extensive hydrolysis. Trivalent ions like Al3+ and Fe3+ are weak acids whose pKa is comparable to that of acetic acid. Solutions of salts such as BeCl2 or Al(NO3)3 in water are noticeably acidic; the hydrolysis can be suppressed by adding an acid such as nitric acid, making the solution more acidic.

Hydrolysis may proceed beyond the first step, often with the formation of polynuclear species. Some "exotic" species such as Sn3(OH)42+ are well characterized. Hydrolysis tends to increase as pH rises leading, in many cases, to the precipitation of an hydroxide such as Al(OH)3 or AlO(OH). These substances, the major constituents of bauxite, are known as laterites and are formed by leaching from rocks of most of the ions other than aluminium and iron and subsequent hydrolysis of the remaining aluminium and iron.

Ions with a formal charge of four have undergone extensive hydrolysis and salts of Zr4+, for example, can only be obtained from strongly acidic solutions. With oxidation states five and higher the concentration of the aqua ion in solution is negligible. In effect, the aqua ion is a strong acid. For example, aqueous solutions of Cr(VI) contain CrO42-.

Cr(H2O)66+ → CrO42- + 2 H2O + 8 H+

Note that reactions such as

Cr2O72- + H2O is in equilibrium with 2 CrO42- + 2 H+

are formally hydrolysis reactions as water molecules are split up yielding hydrogen ions. Such reactions are common among polyoxometalates.